OpenStudy (anonymous):

Christine's sister Claire will be attending college in the fall. In the scale drawing, the dorm room has a length of 4.5 inches. What is the actual length, in feet, of the dorm room?

OpenStudy (anonymous):

where's the drawing? you need the scale in order to do the problem...

OpenStudy (anonymous):

since the scale is 1/4 inch = 1 foot, and the dorm room length on the drawing is 4.5 inches, use the proportion: \(\large \frac{1/4inches}{12inches}=\frac{4.5inches}{x\space inches} \) you'll prolly then need to change your answer to feet instead of inches.